Biashara Limited is newly listed company at the Nairobi Securities Exchange. Last year the company had a successful initial public offer (IPO) and raised $10m. The company however requires additional funds for new investment projects already identified. The finance manager is considering raising the additional funds and has the following information:

  1. Issue 10% preferred stock of $5m at the market price of $150 per share.
  2. Issue debentures of $8m at a market price of $250 per stock debt. The interest to be paid is 10% per annum.
  3. The share capital was issued at $100 each with a floatation cost of $20 per share.
  4. The expected ordinary dividend is $5 growing annually at a constant rate of 2%.
  5. The corporate tax is 40%.

Calculate the expected weighted average cost of capital for Biashara Limited
Answer #1

WACC = (weight of debt * after tax cost of debt) + (weight of preferred stock * cost of preferred stock) + (weight of common stock * cost of common stock)

market value of debt = $8,000,000

market value of preferred stock = $5,000,000

market value of common stock = $10,000,000

total market value = $8,000,000 + $5,000,000 + $10,000,000 = $23,000,000

weight of debt = market value of debt / total market value

weight of debt = $8,000,000 / $23,000,000 = 0.3478

weight of preferred stock = market value of preferred stock / total market value

weight of preferred stock = $5,000,000 / $23,000,000 = 0.2174

weight of common stock = market value of common stock / total market value

weight of common stock = $10,000,000 / $23,000,000 = 0.4348

After-tax cost of debt = interest rate * (1 - tax rate)

After-tax cost of debt = 10% * (1 - 40%) = 6%.

cost of preferred stock = (annual dividend / issue price per share)

annual dividend = face value * dividend rate = $100 * 10% = $10

cost of preferred stock = $10 / $150 = 6.67%

cost of common stock = (expected dividend / net proceeds per share) + growth rate.

net proceeds per share = $100 - $20 = $80

cost of common stock = ($5 / $80) + 2%

cost of common stock = 8.25%

WACC = (weight of debt * after tax cost of debt) + (weight of preferred stock * cost of preferred stock) + (weight of common stock * cost of common stock)

WACC = (0.3478 * 6%) + (0.2174 * 6.67%) + (0.4348 * 8.25%)

WACC = 7.12%